OpenStudy (anonymous):

Solve the following systems by substitution. Substitute the expression in the second equation into the first equation and solve. (If the system has no solution, enter NONE for each answer.) x-y=-8 y=2x+12

OpenStudy (anonymous):

Input y=2x+12 into the x-y=-8 That equals x-2x-12=-8 Solve for x

OpenStudy (anonymous):

Once you solve for x, input x into one of the equations to solve for y

OpenStudy (anonymous):

Do you get it or do you want me to solve it?

OpenStudy (anonymous):

i dont understand where you got the -12 from

OpenStudy (anonymous):

Well y is negative in x-y=-8 so you have to multiply 2x+12 by -1

OpenStudy (anonymous):

so think of it like x-(2x+12)=-8

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes i did thank u so much

OpenStudy (anonymous):

your welcome